• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 16:01:10 on May 3,2008

    Entry number 236722

    keyword=Target status

    As I start shift, the fan is still not working, and loop 1 is
    at 18.15 K, HPH is at 118 W. The target alarm handler is,
    of course, alarming.

    Data-taking is continuing, and previous target operator (Y.C. Wang)
    explains that the target has been stable in this condition for
    several hours. J-P Chen is due in at 16:30 or so, and will make an
    access to restart the fan.

    In the meantime, I step the loop 1 JT valve closed 1 step (why waste
    so much cryogen?). It is now at 38% open. I may gradually step
    the JT closed a few more steps in the meantime.